OpenStudy (anonymous):

Which of the following is not a way to represent the solution of the inequality 8x - (5x + 4) >= -31? x >= -9 -9 <= x A number line with a closed circle on -9 and shading to the left. A number line with a closed circle on -9 and shading to the right

OpenStudy (anonymous):

please help

OpenStudy (aravindg):

k lets simplify (5x + 4) >= -31 5x>=-31-4 5x>=-35 x>=-7

OpenStudy (anonymous):

ohkay

OpenStudy (anonymous):

A number line with a closed circle on -9 and shading to the left.

OpenStudy (anonymous):

explain

OpenStudy (anonymous):

okay

OpenStudy (aravindg):

well look at the inequality we gt ,graphically it is 

OpenStudy (aravindg):

so shading to left is wrong!!

OpenStudy (anonymous):

and that is a closed circle right ?

OpenStudy (aravindg):

yep !

OpenStudy (anonymous):

ohkay, so how do we get to the right answer ?

OpenStudy (anonymous):

8x - (5x + 4) >= -31 8x-5x >= -27 3x >= -27 x >= -9

OpenStudy (aravindg):

oops i didnt see the 8x above srry

OpenStudy (aravindg):

bt still the graph towards right

OpenStudy (anonymous):

sooo... ???

OpenStudy (anonymous):